OpenStudy (anonymous):

Consider the pattern 1, 1/1/2, 1/4/3, 1/7/8,..... describe the pattern and the next 3 numbers.

OpenStudy (lgbasallote):

can you draw how those numbers look like?

OpenStudy (anonymous):

it is one half, one and three fourths, and one and seven eights.

OpenStudy (agent47):

Observe that you are adding three each time here: /1/, /1+3/, /1+3+3/

OpenStudy (agent47):

wait what about the 1/s ?

OpenStudy (anonymous):

yes

OpenStudy (agent47):

they're still there right?

OpenStudy (anonymous):

they are fractions

OpenStudy (agent47):

so basically you have: 1, 2/1, 3/4, 8/7

OpenStudy (anonymous):

its 1/2 not 2/1

OpenStudy (agent47):

no it's 1/1/2, which is 2/1

OpenStudy (anonymous):

yeah sorry

OpenStudy (agent47):

can you take a picture of the original question?

OpenStudy (anonymous):

It is the challenge

OpenStudy (anonymous):

now that is : 1/1 , 3/2 , 7/4, 15/8 , .... so denominators are powers of 2 , 1,2,4,8.... and numerators have difference as powers of 2 3-1=2 , 7-3=4, 15- 7 = 8....... got the pattern ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

I got it now

OpenStudy (anonymous):

so were u able to find next 3 numbers?

OpenStudy (agent47):

ohh that's not 1/1/2, that's 1 and 1/2

OpenStudy (agent47):

no wonder i cldn't find the pattern

OpenStudy (anonymous):

I cant find the next 2 #'s

OpenStudy (anonymous):

denominators will be 16,32,64 right ? 1st numerator will be 15+16=31 got that ? 2nd numerator = 31+32= 63 ok? can u find last ?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

so the next numbers will be 31/16 which is 1 15/16 63/32 which is 1 31/32 and what u got last one as ?

OpenStudy (anonymous):

1/127/128

hartnn (hartnn):

did u forget 64 ? it will be 127/64 which is 1 63/64

OpenStudy (anonymous):

nope i got that one

hartnn (hartnn):

so the next 3 numbers u were searching are : 1 15/16 , 1 31/32 , 1 63/64 right ? 1 127/128 will be the 4th. ok?

OpenStudy (anonymous):

yep i got it